OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 2x + y = 8 −x − y = −5 8 + 2x 8 − 2x −8 − 2x −8 + 2x

OpenStudy (solomonzelman):

the best method here is Elimination. Add the 2 equations together

OpenStudy (solomonzelman):

Ok, they, though, need substitution. Isolate the "y" in the first equation (can you do that?)

OpenStudy (solomonzelman):

2x + y = 8 2x + y - 2x = 8 -2x y = ?

OpenStudy (anonymous):

y = 2

OpenStudy (legends):

Did you get the answer?